OpenStudy (anonymous):

What is the area of this regular polygon? A. 62.5 cm2 B. 148.5 cm2 C. 1075 cm2 D. 2150 cm2

OpenStudy (anonymous):

@ajanijones @iGreen

OpenStudy (igreen):

Picture?

OpenStudy (anonymous):

oh yeah how do I put that?

OpenStudy (anonymous):

Im sorry

OpenStudy (anonymous):

im getting the pic

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

sorry bout that

OpenStudy (igreen):

That is a pentagon. We can use the formula for the area of a pentagon: \(A = \dfrac{5}{2}sa\) Where 's' is the side length and 'a' is the apothem. Plug them in: \(A = \dfrac{5}{2}(25)(17.2)\) Simplify that

OpenStudy (igreen):

Simplify this: \(\dfrac{5 \times 25 \times 17.2}{2}\)

OpenStudy (igreen):

First multiply 5 * 25 * 17.2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2150 cm2

OpenStudy (anonymous):

So I think its 1072 cm2?

OpenStudy (igreen):

Sorry, OS was lagging.

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Now divide by 2.

OpenStudy (igreen):

No, check again.

OpenStudy (anonymous):

I need help with 2 more questions to please?

OpenStudy (anonymous):

I diided it and got 1072 cm2

OpenStudy (anonymous):

1075 im sorry

OpenStudy (igreen):

Sure, just close this one and open a new one.

OpenStudy (igreen):

Yep, that's our answer.

OpenStudy (anonymous):

awsome